PepsiPFU Commissioning Observing - 20140109 UT

Observers: JStorm, IIlyin, MWeber, AJarvinen (LBTO), JHill, DMiller (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: JLittle (LBTO)
Instrument Support:: JMorris (LBTO)
Engineering Support: DAshby (Tucson)

Summary

It seemed like a night filled with minor problems one after the other.

The DX Guide CCD in PEPSIPFU is noisy with faint bars, so we had more problems than you would expect guiding on an 11th magnitude guide star. Jesper measured the photometry, and reports that PFU with Beam Splitter #1 has a delta zeropoint of 5.2 magnitudes compared to a regular Potsdam AGW. This means that 11th magnitude is indeed a fairly faint star.

Then we seemed not to be guiding on the hotspot (both sides). That turned out to be because we were erroneously sending a science wavelength of 1.4 microns from IRAF LBTtools. This looked like a systematic guiding error at low elevation, but was functioning correctly.

DX active optics were seriously running away in Z9 and Z10 even though they worked fine last night. After lots of head scratching Doug cured the problem by power cycling the DX secondary. Apparently there was a peculiar high order shape on the shell that was not removed by reflattening (2 times). This was making a bogus Z9/Z10 signal that caused the divergence of active optics. The recognizable symptom is that some of the S-H spots were significantly extended due to high local slopes.

Doug's zernike_collect program was incorrectly reading a rotator angle, so IDL wasn't converging correctly. And the thresholding (of total wavefront rms) isn't detecting good collimation...... He fixed both of those in real time.

And then there a few human glitches on top of that - decongestants only go so far.

While there were lots of problems, the PEPSI guys did get a significant number of tests completed including setting the ADCs at 33deg elevation.

Details

Sky is clear at twilight.

PCS/PSF preference offloading is enabled.

01:10 Initial collimation struggling because of bright sky and poor seeing.

01:20 Why does left star look OK, but right star looks 2 mags too faint? Pointing is confirmed on a second star.

01:30 Rest and re-set the DX shell to see if that contributed to the funny looking image. It looked the same afterward.

02:24 Move to a brighter star at EL=48. Both guiders agree on the brightness of the star within 10%.

02:30 rollback GCS from GCS.2013D.Ver-2.14.14.0 to GCS.2013D.Ver-2.14.12.0

Restart GCSR, but not yet GCSL.

02:32 ACTIVE preset on right succeeds, but still NOT guiding on the hotspot as indicated by the blue circle or the tcslog.

02:36 Clear Active Optics on Right side.

Doing ADC experiments on Left side. Seeing is 1 arcsec.

02:39 Disable preference offloading.

The Active Optics problem on the right is that Z9 and Z10 are running away and bending the primary into a nasty shape.

03:07:10 Apply -600 Z9 to DX primary (wavefront stable for several images before and after this)

03:17:10 Apply -500 Z10 to DX primary

2014-01-09T03:11:05.897172+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: 10.182, 19.631 [pix] binned
2014-01-09T03:11:05.897184+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: 10.182, 19.631 [pix] unbinned
2014-01-09T03:11:05.897190+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: -0.331, -0.638 [mm] SFP
2014-01-09T03:11:05.897277+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: 1.198 within  limit: 2.080 [arcs] <--- star close enough, ACCEPT
2014-01-09T03:11:05.897405+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: -0.331, -0.638 [mm] SFP
2014-01-09T03:11:05.897417+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: 0.241, -0.677 [mm] SFP (derotated)
2014-01-09T03:11:05.897459+00:00 tcs4 LBT_left_GCS: GT: GS_to_HS offset: 0.401, -1.128 [arcs] SFP
2014-01-09T03:11:05.897468+00:00 tcs4 LBT_left_GCS: GT: GS abs.coord where it should be: 0.004, 0.331 [mm] SFP
2014-01-09T03:11:05.897508+00:00 tcs4 LBT_left_GCS: GT:  pre-starPos update -0.095153, -0.497700 (X,Y) [mm] SFP pre pcsClient.applyGuideCorrection
2014-01-09T03:11:05.897543+00:00 tcs4 LBT_left_GCS: GT: post-starPos update -0.236873, -0.346033 (X,Y) [mm] SFP pre pcsClient.applyGuideCorrection

03:48 collimating on right with Z9 and Z10 turned off (zero gain)

03:51 zernike_collect on right at EL=32.5

Error in gcs_take_image err = AzCam server timed out waiting for full WFSCam image err = WFS image invalid!

Was IRAF sending wavelength of 1.4 microns and overriding PCSInstrument.conf? Yes, that was the guiding "problem" we were seeing at low EL.

03:58 Enable preference offloading again

04:05 SX hexapod times out???

04:10 Huge ACQ offset sent to optics with star only ~2 arcsec from hotspot

04:12 Restart PCS (which disables preference offloading) and GCSL and GCSR (both running v2.14.14)

hexapod timeout is a bit too short

04:49 Doug cycles power on DX M2, and we clear Active Optics on primary again. And that made the DX wavefront look SOOOOOO much better! We think there was a high order shape on the secondary that was masquerading as Z9/Z10.

Measurement of Zernike Rotation Angles on Both Sides

Data in /home/lbto/data/20140109_pfu_left and */home/lbto/data/20140109_pfu_right*'

05:00 John running zernike_collect on the left, and Doug running it on the right with 5-10.dat.

On Left:
     => Zernikes sent to PSF : rms = 317 nm
% Compiled module: IRC_SENDWAVEFRONT.
   Acquiring LEFT WFSC image (exp=15.0000 s)...
     file = left_wfsc000002_dd.fits
      => Zernikes sent to PSF : rms = 176 nm
   Acquiring LEFT WFSC image (exp=15.0000 s)...
     file = left_wfsc000003_dd.fits
      => Zernikes sent to PSF : rms = 225 nm
   Acquiring LEFT WFSC image (exp=15.0000 s)...
     file = left_wfsc000004_dd.fits
      => Zernikes sent to PSF : rms = 298 nm
   Acquiring LEFT WFSC image (exp=15.0000 s)...
     file = left_wfsc000005_dd.fits
      => Zernikes sent to PSF : rms = 364 nm

Ilya may be tuning left hotspot while we are taking data.

Coma runs away with IDL collimation because of a wrong rotator position read by IDL.

05:10 clear active optics everywhere

05:15 zernike_collect again

zernike_collect doesn't recognize convergence, because it is also looking at high order terms.

A good hotspot for the left is 366.2 376.3 for both pupil and fiber position at EL=51deg

Seeing degrades from 1 to 2 arcsec.

05:50 zernike_collect, start=6, limit=600 (with corrected convergence test which makes it much faster)

Save the PFU Pointing

*** Pointing Model Terms ***
IA	-393.99510
IE	 -15.00000
NPAE	  92.82180
CA	 -58.00000
AW	  -3.65500
AN	  23.83830
TX	  50.03400
TF	-124.15230
A1A	   8.10120
IA	-312.16060
IE	  27.00000
NPAE	  53.41900
AW	   6.12880
AN	  23.24110
CA	-184.00000
PZZ1	  15.85850

Guide alignment angle test on DX and then SX

authorized as PFU on both sides as we have been all night

block DX guiding

06:03:12 ptincrement 1 IE 10

06:03:55 ptincrement 1 IE -20

06:04:55 ptincrement 1 IE 10

restore guiding momentarily

06:05:49 ptincrement 1 CA 10

06:06:29 ptincrement 1 CA -20

06:07:08 ptincrement 1 CA 10

restore DX guiding

block SX guiding

06:08:32 ptincrement 0 IE 10

06:09:12 ptincrement 0 IE -18

06:09:39 ptincrement 0 IE 8

restore guiding momentarily

06:10:44 ptincrement 0 CA 8

06:11:22 ptincrement 0 CA -16

06:11:44 ptincrement 0 CA 8

Finish

06:13 Handover to OSURC due to impending moonset. They are switching to LBCs as thin clouds move in.

23:11:50 ptincrement 0 CA 8

Guide image numbers

  • block guide 1291
  • IE +10 1295-1302
  • IE -10 1305-1311
  • IE 0 1312
  • Guide on 1315
  • block guide 1319
  • CA +10 1322-1328
  • CA -10 1331-1337
  • CA 0 1339
  • Guide on 1343-1344

LEFT

  • block guide 1473
  • IE +10 1477-1484
  • IE -8 1486-1493
  • IE 0 1498
  • Guide on 1501-1503
  • Guide off 1507
  • CA +8 1512-1520
  • CA -8 1524-1527
  • CA 0 1530
  • Guide on 1534

-- JohnHill - 09 Jan 2014
Topic revision: r6 - 21 Jan 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback